Exercise 3-6A Identifying transaction type, its effect 011 the accounting equation, and whether
the effect is recorded with a debit or credit


Required
Id ent ify whether each of the follow ing tra nsactio ns is an asse t so urce (AS). asset
use (AU) , asset exchange (A E), or claims exchange (C E). Also e xplai n how each event
afTects th e acco untin g eq uatio n by placing a + for increase, - for decrease. a nd NA for
1101 affec1ed under each of the co mpo nent s of the acco unt in g eq uatio n. Fina lly, in
dicate whet her the efTec t req uires a debit or cred it ent ry. The first event is reco rded as
an examp le.

Stockholders Equity

Event

a


Type of
Event
AS

Assets
+ Debit

Liabilities +
NA


Common Stock
NA


Retained
+ Earnings
+ Credit

a. P rov id ed se rvices on acco un t.
b. Rece ived cash in payment of accounts receivable.
c. Pur c hased la nd by paying cas h.
d. Recognized revenue for services co mple ted: cash collected previo usly.
e. Paid a cash dividend to t he stock holde rs.
f. Paid cash in ad vance for o ne year's rent.
g. Received cash for services lo be performed in the future.
h. 1ncurred other o perat in g expenses o n acco unt.
i. Paid sala rie s pa yable.
j. Recognized expense for prepaid rent that had been used up by the end of th e accounting
period.
k. Provided service s for cas h.
I. P urchased supplie s on accoun t..
m. Recognized expense for supplies used d ur ing the period.
n. Recog nized accrued int erest revenue.
Answer #1

Answer-

Event.   Type of Event   Assets = Liabilities Stockholders’ Equity
+ Common Retained
Stock Earnings
a. AS +Debit = NA + NA + +Credit AQ
b. AE

+Debit

-credit

= NA + NA + NA
c. AE

+Debit

-Credit

= NA + NA + NA
d. CE NA = -Debit + NA + +Credit
e. AU -Credit = NA + NA + -Debit
f. AE

+Debit

-Credit

= NA + NA + NA
g. AS + Debit = +Credit + NA + NA
h. CE NA = +Credit + NA + -Debit
i. AU -Credit = -Debit + NA + NA
j. AU -Debit = NA + NA + -Debit
k. AS +Debit = NA + NA + +Credit
l. AS +Debit = +Credit + NA + NA
m. AU -Credit = NA + NA + -Debit
n. AS +Debit = NA + NA + +Credit